Keep the Highest: /12 Attention: Due to a bug in Google Chrome, this page may not function correctly.Click here to learn more a Aa 4. Market value ratios Ratios are mostly calculated using data drawn from the financial statements of a firm. However, another group of integrating information from both the market and the firms financial statements Consider the case of Cute Camel Woodcraft Company: atos,caled market baedtos, relate to ftms observable market value, stock prices, and book values, Cute Camel Woodcraft Company just reported earnings current stock price of $34.00 per share. The company is forecasting after tax (also called net income) of $9,750,000, and a an increase of 25% for its after-tax income next year, 5,500,000 to 8,400,000). but it also expects it will have to issue 2,900,000 new shares of stock (raising its shares outstanding from If Cute Camels forecast tums out to be correct and its price-to-earnings (P/E) ratio does not change, what does the companys management expect its stock price to be one year from now? (Round any P/E ratio calculation to four decimal places). $27.85 per share $34.00 per share O $20.89 per share $34.81 per share One year later, Cute Camels shares are trading at $53.32 per share, and the company reports the value of its total common equity as $46,788,000. Given this information, Cute Camelrs market-to-book (M/B) ratio is Can & companys shares exhibit a negative P/E ratio? O Yes Which of the following statements is true about market value ratios? O Companies with high research and development (RAD) expenses tend to have low P/E ratios. O Compenies with high research and development (R&D)expenses tend to have high P/E ratios. O Type here to search PrtScn 4
Answer #1

Answer to Part 1:

P/E Ratio = Price per Share / Earnings per Share

Current Situation:
Earnings per Share = Net Income / Shares Outstanding
Current Earnings per Share = 9,750,000 / 5,500,000
Current Earnings per Share = $1.77

Current P/E Ratio = 34 / 1.77
Current P/E Ratio = 19.21 times

Proposed Situation:
Proposed Net Income = $9,750,000 * 1.25
Proposed Net Income = $12,187,500

Proposed Earnings per Share = 12,187,500 / 8,400,000
Proposed Earnings per Share = $1.45

P/E Ratio = Price per Share / Earnings per Share
19.21 = Price per Share / 1.45
Price per Share = $27.85

Answer to Part 2:

Market to Book Ratio (M/B) = Market Value / Book Value
Market Value = $53.32 * 8,400,000
Market Value = $447,888,000

Market to Book Ratio (M/B) = $447,888,000 / $46,788,000
Market to Book Ratio (M/B) = 9.57

Answer to Part 3:

Yes, as the Company can incur a net loss leading to negative P/E ratio.

Answer to Part 4:

Companies with high research and development (R&D) expenses tend to have high P/E Ratios.